# On the Multiary Algebraic Formulation of an Idempotent Symmetric Limit Convex Structure

### v1

@@ 3.1 Intermediate Points and Copositivity: the Case of two Points
To achieve this, a notion of intermediate points was introduced in [9]. Below, we briefly summarize the key results. To show that $Co^{\infty}(x,y)$ is the Painlevé-Kuratowski limit of the sequence $\{Co^{(p)}(x,y)\}_{p\in\mathbb{N}}$, it is useful to define the concept of $i$-intermediate points of order $p$.
For all $(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, let $\mathcal{I}(x,y)$ be the subset defined by $\mathcal{I}(x,y)=\{i\in[n]:x_{i}y_{i}<0\}$ and let $n(x,y)$ be its cardinal. For all natural numbers $p$, let us consider the map $\gamma^{(p)}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times[0,+\infty]\longrightarrow\mathbb{R}$ defined by:
@@
Remark that $\gamma(x,y,0)=x$. For all $i\in\mathcal{I}(x,y)$ and all $t_{i}^{\star}\in\mathbb{R}_{++}$ a point $\gamma\in\mathbb{R}^{n}$ is called a $i$-intermediate point between $x$ and $y$ if there is some $t_{i}^{\star}\in]0,+\infty[$ such that $\big(\gamma(x,y,t_{i}^{\star})\big)_{i}=0.$ For all $i\in\mathcal{I}(x,y)$, the map $t\mapsto\gamma_{i}(x,y,t)$ has a unique zero $t_{i}^{\star}=-\frac{x_{i}}{y_{i}}=|\frac{x_{i}}{y_{i}}|>0$ and there is a unique $i$-intermediate point
@@
In general given $a_{1},...,a_{m}\in\mathbb{R}$, if $(I_{1},I_{2})$ is a partition of $[m]$ there is no guarantee that $\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in[I_{1}]}a_{i}\Big)\boxplus\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{2}}a_{i}\Big)$ is identical to $\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{1}\cup I_{2}}a_{i}$. However, this is true if $\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{1}}a_{i}$ and $\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{2}}a_{i}$ are not symmetric or equal to 0. Since for all $m$, $\gamma(x,y,t_{i_{m}}^{\star})$ and $\gamma(x,y,t_{i_{m+1}}^{\star})$ are copositive, we have for all $r,s\in[0,1]$ with $\max\{r,s\}=1$ the relation:
@@
Suppose that there exists $m$ subsets of $\mathbb{N}$, $I_{1},I_{2}...I_{m}$, with $x^{(i)}\in\mathbb{R}^{n}$ for all $i\in I_{j}$. Suppose moreover that for all $j,k$ $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{j}}x^{(i)}$ and $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{k}}x^{(i)}$ are copositive i.e. have the same sign which equivalently means $\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{j}}x^{(i)}\Big)\odot\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{k}}x^{(i)}\Big)\geq 0$. Then:
@@ 3.2 Intersection with $n$-dimensional Orthant, Extreme Points in Limits and Inclusion Property
as an intermediate point of order $p$ for $A$, indexed by the pair $(I,J)$ if $t_{{}_{I},_{J}}^{(p)}$ is the unique element of $\Delta_{J}^{(p)}$ satisfying
@@
In the following, given a $\varphi_{p}$-convex subset $C$ of $\mathbb{R}^{n}$, we say that $z$ is a $\varphi_{p}$-extreme point if there do not exist $u,v\in C$ and $(s,t)\in\Delta_{2}^{(p)}$, such that $z=su\stackrel{{\scriptstyle p}}{{+}}tv$. Given a $n$-dimensional orthant $K$, the subset $Co^{p}(A)\cap K$ has a $\varphi_{p}$-internal representation. This means that there exists a finite subset $\mathcal{E}_{K}^{(p)}(A)$ of $\varphi_{p}$-extreme points such that
@@
In the following, we say that an intermediate point is properly indexed on $(I,J)$ if $\Lambda_{{}_{I},_{J}}$ is $\varphi_{p}$-invertible. In the next lemma, we show that any $\varphi_{p}$-extreme point is a an intermediate point of order $p$ properly indexed on some pair $(I,J)$. In the following, $\zeta^{(p)}(A)$ denotes the set of all the intermediate points of order $p$ for $A$ that are properly indexed.
@@
Proof: Suppose that there exists some $j$ such that $z=x^{(j)}$. Then, by definition, $z$ is an intermediate point of order $p$ properly indexed on the pair $\emptyset,\{j\}$}, and the result follows immediately.
@@
Now, suppose that $\mathrm{Card},I<n_{J}-1$, with $z_{i}\neq 0$ for $i\notin I$. We will derive a contradiction. Let $H$ be a $\varphi_{p}$-hyperplane of dimension $n_{J}-1$ passing through $x^{(j)}$ for all $j\in J$. Since $t_{j}>0$, there exists a ball $B_{\infty}(z,r]$ for some $r>0$ such that
@@
One can see form equation (49) that each term of the $\varphi_{p}$-generalized sum is independent of $p$. The limit of $\big|\Lambda_{{}_{I},_{J}}\big|_{p}$ being computed independently. An important property established in [12] is that for all $x\in\mathbb{R}^{n}$, we have
@@
The next result is then immediate.
@@
$x\in\Big\{\operatorname*{\text{\Huge${}_{\boxplus}$}}_{(I,J)\in\mathcal{C}}t_{{}_{{}_{I},_{J}}}\zeta_{{}_{I},_{J}}:\max_{{}_{(I,J)}\in\mathcal{C}}t_{{}_{{}_{I},_{J}}}=1,t_{{}_{{}_{I},_{J}}}\geq 0\Big\}.\quad\Box$
@@
Let us denote $\ell_{I}=(n_{I}+1)!$ and consider a bijective index map $i:[\ell_{I}]\longrightarrow\mathfrak{S}_{n_{J}}$ where $n_{I}=\operatorname*{\mathrm{Card}}\,I$, which associates to any number in $\ell_{I}=(n_{I}+1)!$ a permutation $\sigma\in\mathfrak{S}_{n_{J}}$. For all $j\in J$, and all $q\in[\ell_{I}]$, let us denote $\alpha_{\sigma,j}=s_{i(\sigma),j}=s_{q,j}$. By construction, it follows that
@@
Hence there is some $I_{0},J_{0},q_{0},j_{0}$ such that $r_{I_{0},J_{0}}=1$ and $s_{k_{0},j_{0}}=1$. Thus:
@@
In subsection 3.4 we give an example of decomposition of a $\mathbb{B}$-polytope in a three-dimensional space in section. This example is depicted in Figure 3.2.2.
@@ 3.3 Converse Inclusion and Algebraic Formulation of a Polytope
It was shown in [11] that for all finite non-empty subsets $A$ of $\mathbb{R}^{n}$, the sequence $\{Co^{p}(A)\}_{p\in\mathbb{N}}$ converges to $Co^{\infty}(A)$ with respect to the Hausdorff metric. Proposition 3.3.2 provides an equivalent formulation of this result, since $Co^{\infty}(A)$ is compact.
In what follows, we will show that this definition is equivalent to the case where we consider the convex hull of two points.
@@
A subset $C$ of $\mathbb{R}^{n}$ is $\mathbb{B}$-convex if for all finite subsets $A=\{x^{(1)},...,x^{(m)}\}$ of $\mathbb{R}^{n}$
@@
A subset $C$ of $\mathbb{R}^{n}$ is $\mathbb{B}$ if for any finite subset $A=\{x^{(1)},\dots,x^{(m)}\}$ of $C$ and for all $t\in\prod_{j\in[m]}\mathbb{R}^{n_{j}}$ with $n_{j}\in\mathbb{N}$ for each $j$, the conditions
@@
The advantage of this formulation is that it is sufficiently general to extend the class of functions defined in [1]. Likewise, it would be interesting to explore a harmonic version of this convexity structure in order to generalize certain results established in [2]. Note that the convexity formalism introduced in this paper is consistent with the concept of idempotent symmetric spaces defined in [13]. It is easy to see that an idempotent symmetric space is $\mathbb{B}$-convex. Notice also that a non-associative symmetrization of Max-Plus convex sets was proposed in [12, 13], and similar investigations could be carried out. This is briefly in the following example.
@@ 3.4 An Example and some Implications
In this section, we present a detailed example illustrating the construction of a 3-dimensional simplex. Specifically, we provide a counterexample demonstrating that for any finite part $A$, $\mathrm{IS}[A]\nsubseteq Co^{\infty}(A)$. In other words, a $\mathbb{B}$-polytope is not idempotent symmetric convex, i.e. it does not satisfy the condition of convexity at the limit for two points, as defined in Definition [9] (which represents a special case of $\mathbb{B}$-convexity).
It was establish in [11] that if $A=\{x^{(1)},...,x^{(m)}\}$ is a finite subset of $\mathbb{R}^{n}$, then for all natural numbers $p$, and all $K\in\mathcal{K}(A)$, since $\mathcal{Z}^{(p)}(A)\cap K$ contains all the $\varphi_{p}$-extreme points of $Co^{p}(A)\cap K$, that we have $Co^{p}\big(\mathcal{Z}^{(p)}(A)\cap K\big)=Co^{p}(A)\cap K.$ Moreover
@@
However, it is important to emphasize that the proposed formulation, which involves any arbitrary finite subset of the considered set, provides numerous advantages. Firstly, it is defined in algebraic terms and is very useful for solving the separation problem, as shown in the next section. Moreover, it explicitly provides a polytope structure that, while not coinciding with the convex hull, is nonetheless the limit of a sequence of $\varphi_{p}$-convex hulls.
@@
Proof: To prove this result, we need to show that that for all finite subset $A=\{x^{(1)},...,x^{(m)}\}\subset\mathbb{R}^{2}$ we have:
@@
4 From the Seprataion of Polytope to the Separation of Convex Set
@@ 4.1 Limit of Idempotent Lower (Upper) Symmetric Halfspaces
Let $\langle\cdot,\cdot\rangle_{p}$ denote the $\varphi_{p}$-inner product defined for any $p\in\mathbb{N}$ and all $x,y\in\mathbb{R}^{n}$ as $\langle x,y\rangle_{p}=\varphi_{p}^{-1}\big(\langle\phi_{p}(x),\phi_{p}(y)\rangle\big)$ (see [9, 10]). In the following we introduce the operation $\langle\cdot,\cdot\rangle_{\infty}:\mathbb{R}^{n}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}$ defined for all $x,y\in\mathbb{R}^{n}$ by $\langle x,y\rangle_{\infty}=\lim_{p\longrightarrow\infty}\langle x,y\rangle_{p}=\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in[n]}x_{i}y_{i}$. Let $\|\cdot\|_{\infty}$ be the Tchebychev norm defined by $\|x\|_{\infty}=\max_{i\in[n]}|x_{i}|$. It is established in [9] that for all $x,y\in\mathbb{R}^{n}$, we have: $(i)$ $\sqrt{\langle x,x\rangle_{\infty}}=\|x\|_{\infty}$; $(ii)$ $|\langle x,y\rangle_{\infty}|\leq\|x\|_{\infty}\|y\|_{\infty}$; $(iii)$ For all $\alpha\in\mathbb{R}$, $\alpha\langle x,y\rangle_{\infty}=\langle\alpha x,y\rangle_{\infty}=\langle x,\alpha y\rangle_{\infty}$.
@@
In [10], the geometric deformation of a sequence of $\varphi_{p}$-halfspaces, expressed as $H_{p}=[f_{p}\leq c_{p}]$, was analyzed. For each natural number $p$, $f_{p}:\mathbb{R}^{n}\longrightarrow\mathbb{R}$ is a $\varphi_{p}$-linear form defined by $f(x)=\langle a^{(p)},x\rangle_{p}$, with $(a^{(p)},c_{p})\in\mathbb{R}^{n}\times\mathbb{R}$. It was shown that the Kuratowski-Painlevé limit of this sequence can be expressed with respect to the idempotent symmetric forms.
@@
The subsequent result is closely related to Proposition 4.1.1. Although the latter could, in principle, be employed to establish it via Hausdorff-type arguments, we present here a fully self-contained proof. Such an approach will prove instrumental in extending our analysis from the separation of polytopes to the more general framework of compact $\mathbb{B}$-convex sets.

### v2

@@ 3.1 Intermediate Points and Copositivity: the Case of two Points
To achieve this, a notion of intermediate points was introduced in [9]. Below, we briefly summarize the key results. To show that $Co^{\infty}(x,y)$ is the Painlevé-Kuratowski limit of the sequence $\{Co^{(p)}(x,y)\}_{p\in\mathbb{N}}$, it is useful to define the concept of $i$-intermediate points of order $p$, where $i\in[n]$ is such that $x_{i}y_{i}<0$.
For all $(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, let $\mathcal{I}(x,y)$ be the subset defined by $\mathcal{I}(x,y)=\{i\in[n]:x_{i}y_{i}<0\}$ and let $n(x,y)$ be its cardinal. For any natural number $p$, let us consider the map $\gamma^{(p)}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times[0,+\infty]\longrightarrow\mathbb{R}$ defined by:
@@
Remark that $\gamma(x,y,0)=x$. For all $i\in\mathcal{I}(x,y)$ and all $t_{i}^{\star}\in\mathbb{R}_{++}$, a point $\gamma\in\mathbb{R}^{n}$ is called an $i$-intermediate point between $x$ and $y$ if there is some $t_{i}^{\star}\in]0,+\infty[$ such that $\big(\gamma(x,y,t_{i}^{\star})\big)_{i}=0.$ For all $i\in\mathcal{I}(x,y)$, the map $t\mapsto\gamma_{i}(x,y,t)$ has a unique zero $t_{i}^{\star}=-\frac{x_{i}}{y_{i}}=|\frac{x_{i}}{y_{i}}|>0$ and there is a unique $i$-intermediate point
@@
In general given $a_{1},...,a_{m}\in\mathbb{R}$, if $(I_{1},I_{2})$ is a partition of $[m]$ there is no guarantee that $\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{1}}a_{i}\Big)\boxplus\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{2}}a_{i}\Big)$ is identical to $\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{1}\cup I_{2}}a_{i}$. However, this is true if $\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{1}}a_{i}$ and $\operatorname*{\text{\Huge${}_{\boxplus}$}}\limits_{i\in I_{2}}a_{i}$ are not symmetric or are equal to 0. Since for all $m$, $\gamma(x,y,t_{i_{m}}^{\star})$ and $\gamma(x,y,t_{i_{m+1}}^{\star})$ are copositive, we have for all $r,s\in[0,1]$ with $\max\{r,s\}=1$ the relation:
@@
Suppose that there exists $m$ subsets of $\mathbb{N}$, $I_{1},I_{2}...I_{m}$, with $x^{(i)}\in\mathbb{R}^{n}$ for all $i\in I_{j}$. Suppose moreover that for all $j,k$, $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{j}}x^{(i)}$ and $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{k}}x^{(i)}$ are copositive i.e. have the same sign which equivalently means $\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{j}}x^{(i)}\Big)\odot\Big(\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in I_{k}}x^{(i)}\Big)\geq 0$. Then:
@@ 3.2 Intersection with $n$-dimensional Orthant, Extreme Points in Limits and Inclusion Property
as an intermediate point of order $p$ for $A$, indexed by the pair $(I,J)$, if $t_{{}_{I},_{J}}^{(p)}$ is the unique element of $\Delta_{J}^{(p)}$ satisfying
@@
In the following, given a $\varphi_{p}$-convex subset $C$ of $\mathbb{R}^{n}$, we say that $z$ is a $\varphi_{p}$-extreme point if there do not exist $u,v\in C$ and $(s,t)\in\Delta_{2}^{(p)}$, such that $z=su\stackrel{{\scriptstyle p}}{{+}}tv$. Given an $n$-dimensional orthant $K$, the subset $Co^{p}(A)\cap K$ has a $\varphi_{p}$-internal representation. This means that there exists a finite subset $\mathcal{E}_{K}^{(p)}(A)$ of $\varphi_{p}$-extreme points such that
@@
In the following, we say that an intermediate point is properly indexed on $(I,J)$ if $\Lambda_{{}_{I},_{J}}$ is $\varphi_{p}$-invertible. In the next lemma, we show that any $\varphi_{p}$-extreme point is a an intermediate point of order $p$ properly indexed on some pair $(I,J)$. In the following, $\mathcal{Z}^{(p)}(A)$ denotes the set of all the intermediate points of order $p$ for $A$ that are properly indexed.
@@
Proof: Suppose that there exists some $j$ such that $z=x^{(j)}$. Then, by definition, $z$ is an intermediate point of order $p$ properly indexed on the pair $({\emptyset,\{j\}})$, and the result follows immediately.
@@
Now, suppose that $\mathrm{Card}\,I<n_{J}-1$, with $z_{i}\neq 0$ for $i\notin I$. We will derive a contradiction. Let $H$ be a $\varphi_{p}$-hyperplane of dimension $n_{J}-1$ passing through $x^{(j)}$ for all $j\in J$. Since $t_{j}>0$, there exists a ball $B_{\infty}(z,r]$ for some $r>0$ such that
@@
One can see from equation (49) that each term of the $\varphi_{p}$-generalized sum is independent of $p$, the limit of $\big|\Lambda_{{}_{I},_{J}}\big|_{p}$ being computed independently. An important property established in [12] is that for all $x\in\mathbb{R}^{n}$, we have:
@@
The next result is quite immediate.
@@
$x\in\Big\{\operatorname*{\text{\Huge${}_{\boxplus}$}}_{(I,J)\in\mathcal{C}}t_{{}_{{}_{I},_{J}}}\zeta_{{}_{I},_{J}}:\max_{{}_{(I,J)}\in\mathcal{C}}t_{{}_{{}_{I},_{J}}}=1,t_{{}_{{}_{I},_{J}}}\geq 0\Big\}.$
Since for any $x\in\operatorname*{\mathrm{Ls}}_{p\longrightarrow\infty}Co^{p}(A)$ such a collection $\mathcal{C}$ exists and since the set of all the possible pairs $(I,J)$ is finite, the result follows immediately. $\Box$
@@
Let us denote $\ell_{I}=(n_{I}+1)!$ and consider a bijective index map $i:[\ell_{I}]\longrightarrow\mathfrak{S}_{n_{J}}$ where $n_{I}=\operatorname*{\mathrm{Card}}\,I$, which associates to any number in $\ell_{I}]$ a permutation $\sigma\in\mathfrak{S}_{n_{J}}$. For all $j\in J$, and all $q\in[\ell_{I}]$, let us denote $\alpha_{\sigma,j}=s_{i(\sigma),j}=s_{q,j}$. By construction, it follows that
@@
Hence there is some $I_{0},J_{0},q_{0},j_{0}$ such that $r_{{}_{I_{0},J_{0}}}=1$ and $s_{k_{0},j_{0}}=1$. Thus:
@@
In subsection 3.4 we give an example of decomposition of a $\mathbb{B}$-polytope in a three-dimensional space. This example is depicted in Figure 3.2.2.
@@ 3.3 Converse Inclusion and Algebraic Formulation of a Polytope
It was shown in [11] that for all finite non-empty subsets $A$ of $\mathbb{R}^{n}$, the sequence $\{Co^{p}(A)\}_{p\in\mathbb{N}}$ converges to $Co^{\infty}(A)$ with respect to the Hausdorff-Pompeiu metric. Proposition 3.3.2 provides an equivalent formulation of this result, since $Co^{\infty}(A)$ is compact.
In what follows, we show that, in the case of two points, this formula is equivalent to the one obtained in [9].
@@
A subset $C$ of $\mathbb{R}^{n}$ is $\mathbb{B}$-convex if for all finite subsets $A=\{x^{(1)},...,x^{(m)}\}$ of $C$
@@
A subset $C$ of $\mathbb{R}^{n}$ is $\mathbb{B}$-convex if for any finite subset $A=\{x^{(1)},\dots,x^{(m)}\}$ of $C$ and for all $t\in\prod_{j\in[m]}\mathbb{R}^{n_{j}}$ with $n_{j}\in\mathbb{N}$ for each $j$, the conditions
@@
The advantage of this formulation is that it is sufficiently general to extend the class of functions defined in [1]. Likewise, it would be interesting to explore a harmonic version of this convexity structure in order to generalize certain results established in [2]. Note that the convexity formalism introduced in this paper is consistent with the concept of idempotent symmetric pseudo-fields and idempotent symmetric spaces defined in [13]. It is easy to see that an idempotent symmetric space is $\mathbb{B}$-convex. Notice also that a non-associative symmetrization of Max-Plus convex sets was proposed in [12, 13], and similar investigations could be carried out. This is briefly done in the following example.
@@ 3.4 An Example and some Implications
In this section, we present a detailed example illustrating the construction of a 3-dimensional simplex. Specifically, we provide a counterexample demonstrating that for any finite part $A$, $\mathrm{IS}[A]\nsubseteq Co^{\infty}(A)$. In other words, in general, a $\mathbb{B}$-polytope is not idempotent symmetric convex, i.e. it does not satisfy the condition of convexity at the limit for two points, as defined in [9] (which represents a special case of $\mathbb{B}$-convexity).
It was established in [11] that if $A=\{x^{(1)},...,x^{(m)}\}$ is a finite subset of $\mathbb{R}^{n}$, then for all natural numbers $p$, and all $K\in\mathcal{K}(A)$, since $\mathcal{Z}^{(p)}(A)\cap K$ contains all the $\varphi_{p}$-extreme points of $Co^{p}(A)\cap K$, that we have $Co^{p}\big(\mathcal{Z}^{(p)}(A)\cap K\big)=Co^{p}(A)\cap K.$ Moreover
@@
However, it is important to emphasize that the proposed formulation, which involves any arbitrary finite subset of the considered set, provides numerous advantages. Firstly, it is also defined in algebraic terms and is very useful for solving the separation problem, as shown in the next section. Moreover, it explicitly provides a polytope structure that, while not coinciding with the convex hull, is nonetheless the limit of a sequence of $\varphi_{p}$-convex hulls.
@@
Proof: To prove this result, we need to show that for all finite subsets $A=\{x^{(1)},...,x^{(m)}\}\subset\mathbb{R}^{2}$ we have:
@@
4 From the Separation of Polytopes to the Separation of Convex Sets
@@ 4.1 Limit of Idempotent Lower (Upper) Symmetric Halfspaces
Let $\langle\cdot,\cdot\rangle_{p}$ denote the $\varphi_{p}$-inner product defined for any $p\in\mathbb{N}$ and all $x,y\in\mathbb{R}^{n}$ as $\langle x,y\rangle_{p}=\varphi_{p}^{-1}\big(\langle\phi_{p}(x),\phi_{p}(y)\rangle\big)$ (see [9, 10]). In the following we introduce the operation $\langle\cdot,\cdot\rangle_{\infty}:\mathbb{R}^{n}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}$ defined for all $x,y\in\mathbb{R}^{n}$ by $\langle x,y\rangle_{\infty}=\lim_{p\longrightarrow\infty}\langle x,y\rangle_{p}=\operatorname*{\text{\Huge${}_{\boxplus}$}}_{i\in[n]}x_{i}y_{i}$. Let $\|\cdot\|_{\infty}$ be the Tchebychev norm defined by $\|x\|_{\infty}=\max_{i\in[n]}|x_{i}|$. It was established in [9] that for all $x,y\in\mathbb{R}^{n}$, we have: $(i)$ $\sqrt{\langle x,x\rangle_{\infty}}=\|x\|_{\infty}$; $(ii)$ $|\langle x,y\rangle_{\infty}|\leq\|x\|_{\infty}\|y\|_{\infty}$; $(iii)$ For all $\alpha\in\mathbb{R}$, $\alpha\langle x,y\rangle_{\infty}=\langle\alpha x,y\rangle_{\infty}=\langle x,\alpha y\rangle_{\infty}$.
@@
In [10], the geometric deformation of a sequence of $\varphi_{p}$-halfspaces, expressed as $H_{p}=[f_{p}\leq c_{p}]$, was analyzed. For each natural number $p$, $f_{p}:\mathbb{R}^{n}\longrightarrow\mathbb{R}$ is a $\varphi_{p}$-linear form defined by $f(x)=\langle a^{(p)},x\rangle_{p}$, with $(a^{(p)},c_{p})\in\mathbb{R}^{n}\times\mathbb{R}$. It was shown that the Painlevé-Kuratowski limit of this sequence can be expressed with respect to the idempotent symmetric forms.
@@
The subsequent result is closely related to Proposition 4.1.1. Although the latter could, in principle, be employed to establish it via Hausdorff-Pompeiu type arguments, we present here a fully self-contained proof. Such an approach will prove instrumental in extending our analysis from the separation of polytopes to the more general framework of compact $\mathbb{B}$-convex sets.